OpenStudy (anonymous):

Complex Numbers Number of values of z (real or complex) simultaneously satisfying the system of equations 1+z+z^2 .... z^17 = 0 and 1+z+z^2 .... z^13 = 0

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

I'm glad I was thought worthy of this. *bows*\[\dfrac{z^{18} - 1}{z -1}=0\]\[\dfrac{z^{14} -1}{z-1}=0\]As \(z=1\) is not a solution, the problem boils down to:\[z^{18}=1\]\[z^{14} =1 \]

Parth (parthkohli):

So the answer should be gcd(18,14) = 2?

OpenStudy (anonymous):

lemme show my work.. first 14 terms or 1st exp=0 then z^14(1+z+z^2+z^3) = 0 that gives 4 solution z=0 and three from the cubic z=0 is rejected but the answer is 1

Parth (parthkohli):

Oh, hold on.

Parth (parthkohli):

We started with the assumption that z = 1 is not a solution (obviously). z^14 = 1 and z^18 = 1 have solutions 1 and -1 by simple inspection. But 1 is eliminated.

Parth (parthkohli):

So only -1 satisfies this, yes?

OpenStudy (anonymous):

yeah..but whats problem in my solution?

Parth (parthkohli):

Hmm, tricky solution.

OpenStudy (anonymous):

and we are talking about complex number so we can get 14 root of x^14= 1 and 18 of x^18 = 1 using DMT.. so will there be any common root is there is then more roots

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 1+z+z^2+\cdots+z^{17}=0\hspace{.33em}\\~\\ 1+z+z^2+\cdots+z^{13}=0\hspace{.33em}\\~\\ \implies z^{14}+z^{15}+z^{16}+z^{17}=0\hspace{.33em}\\~\\ \implies z^{14}\left(1+z+z^{2}+z^{3}\right)=0\hspace{.33em}\\~\\ \implies 1+z+z^{2}+z^{3}=0\hspace{.33em}\\~\\ \implies 1+z+z^2(1+z)=0\hspace{.33em}\\~\\ \implies (1+z^2)(1+z)=0\hspace{.33em}\\~\\ z=-1,-i,i\hspace{.33em}\\~\\ \end{align}}\)

Parth (parthkohli):

\[z^2(z + 1) + 1(z+1) = (z^2+1)(z+1)\]The reason why \(\pm i \) is rejected is because the sum of any four consecutive powers is zero. But I'd have to come up with a better explanation.

OpenStudy (rational):